# Safe and Optimal Learning from Preferences via Weighted Temporal Logic with Applications in Robotics and Formula 1

### v1

@@ III Problem Statement
Learning from human feedback in autonomous systems broadly encompasses problems such as learning from preferences, demonstrations, or rankings. Classical approaches include inverse reinforcement learning and behavioral cloning, which directly infer policies or reward structures from data [20, 21]. Both methods typically require a reinforcement learning-style framework. In contrast, we propose to learn a utility function that captures the preferences independent of a specific agent or environment. Ideally, this utility function serves as a unifying representation of human intent, and whenever trajectories consistent with these preferences are required, they can be synthesized via downstream trajectory generation methods. To restrict the search space of possible utility functions, we focus on a structured family of functions defined by the safety requirements or task specifications. Particularly, we leverage the expressive power of WSTL to describe specifications and the weighted robustness. By parameterizing weights in the quantitative semantics, we enable different weight valuations to encode potential instantiations of the utility function. The learned WSTL formula can be seamlessly integrated to control synthesis frameworks such as model predictive control (MPC), to generate safe and personalized trajectories. Moreover, as weight valuations represent the importance of subformulas or time instances, the result of the learning framework is interpretable, in contrast to using neural network-like structures for learning.
In other words, learning from human feedback problems via WSTL requires a task description written in STL and a user dataset. It searches for a weighted quantitative semantics that maximizes the problem’s objective. Formally, we define the problem as follows:
Problem 1 (Learning from Human Feedback (LHF)).
@@
$\begin{array}[]{ccc}w^{*}\in&\max\limits_{w\in\mathbb{R}^{n}_{>0}}&f(\phi_{w},\mathcal{D})\\ &\text{ s.t. }&\text{dataset constraints}\end{array}.$ (3)
The form of the objective function and dataset constraints depend on the learning setup and the available data. For safe preference learning, the user dataset is given as $\mathcal{PR}=\{(\sigma_{i_{1}},\sigma_{i_{2}},l_{i})\}_{i=1}^{P}$, where $\sigma_{i_{1}}$ and $\sigma_{i_{2}}$ are compared and $l_{i}\in\{1,-1\}$ denotes its label: $l_{i}=1$, when $\sigma_{i_{1}}$ is preferred over $\sigma_{i_{2}}$ and $l_{i}=-1$ otherwise. The goal is to maximize the number of correctly ordered pairs, leading to the objective function of the form $f(\phi_{w},\mathcal{PR})=\sum_{i=1}^{P}\mathbb{1}_{\big(l_{i}\Delta_{i}(w)>0\big)}$, with constraints $\Delta_{i}(w)=r(\sigma_{i_{1}},\phi_{w})-r(\sigma_{i_{2}},\phi_{w})$. For learning-to-rank, the dataset is an ordered set $\mathcal{R}=\{\sigma_{1}\succeq\sigma_{2}\succeq\ldots\succeq\sigma_{P}\}$. This can be reformulated as a pairwise preference dataset $\mathcal{PR}=\{(\sigma_{i},\sigma_{j},1)\}_{i=1,j=i+1}^{P-1,P}$, reducing the problem to safe preference learning, with the same objective function and $P\choose 2$ constraints. Finally, in learning from demonstrations, we are given a set of “desired” trajectories $\mathcal{D}=\{\sigma_{i}\}_{i=1}^{D}$. In this setup, we can use $f(\phi_{w},\mathcal{D})=\sum_{i=1}^{D}r(\sigma_{i},\phi_{w})$ as the objective value, which aims to assign high weighted robustness values to desired trajectories.
The computational complexity of WSTL synthesis problems grows with the complexity of the formula. In general, computing the weighted robustness for a signal involves evaluating predicate values at multiple time instances, multiplying them with corresponding weight valuations, and applying nested min/max operations, following the WSTL semantics. Since weights appear multiplicatively, optimizing over them results in a mixed integer program (MIP) with multi-linear constraints, where integers appear as in [22]. As such, weight synthesis problems were solved with heuristic approaches such as gradient descent [23] or sampling [10] in the literature.
@@ IV Reduction to MILP
In this paper, we propose two procedures to reduce the problem size and linearize the constraints, making the problem easier to solve. The first procedure, called the structural pruning, systematically examines how values in the Robustness Computation Tree are propagated to the root node to determine the robustness value of a signal with respect to a formula. Based on this observation, we can prune the branches that cannot affect the robustness value regardless of the weight valuations assigned. The second procedure, called the $\log$-transform, leverages the product identity of the logarithm function to convert constraints in the multiplication form into constraints in the summation form. Since the logarithm is only defined for positive values, a naive approach requires that all subformulas (subtrees in RCT) must have positive robustness for the transformation to be valid, which is a restrictive requirement in practice. To address this, we combine the $\log$-transform with structural pruning, and ensure that only contributing parts with the same robustness value sign of the formula are considered in the robustness computation, therefore making all variables of the constraints that encode this computation have the same sign. Together, these two procedures allow the problem to be reformulated as an MILP with a potentially reduced number of constraints.
Before getting into the details of the procedures, we note that all formulas are assumed to be in positive normal form (PNF), that is, negations appear only in front of predicates. While this assumption is not restrictive, as negations can be pushed to the predicate level, we discuss the rationale later.
@@ IV-A Structural Pruning
The key insight of the structural pruning is that if the overall robustness of a signal is positive (indicating satisfaction), then any subformula with negative or zero robustness values must be “absorbed” by one of its parent operations in the Robustness Computation Tree. Consequently, when fitting weights to this tree, we know that weights connecting nodes with negative or zero robustness value to their parents cannot change the robustness. This allows us to systematically prune parts of the tree whose values are not actively contributing to robustness.
@@
1:function Prune($\sigma,\phi,t,s_{\phi}$)
2:if Root node of $\mathcal{T}_{\phi,t}$ is an Operator then
3:  for $(\phi_{i},t^{\prime})\in$ Children of the Root do
4:   if $\mathrm{sign}(\rho(\sigma,\phi_{i},t^{\prime}))=s_{\phi}$ then
5:    $\mathcal{T}_{\phi_{i},t}\leftarrow\textsc{Prune}(\sigma,\phi_{i},t^{\prime},s_{\phi_{i}})$ $\triangleright$ Prune child
6:   else
7:    $\mathcal{T}_{\phi,t}\leftarrow\mathcal{T}_{\phi,t}\setminus\mathcal{T}_{\phi_{i},t^{\prime}}$ $\triangleright$ Delete child
8:   end if
9:  end for
[…]
11:  return $\mathcal{T}_{\phi,t}$ $\triangleright$ Return the tree
12:end if
13:end function
@@
Example 1 (continued).
Consider $\phi$, and $\sigma=[5,6,7,-1,4,2]$. Figure 2 shows the pruned RCT, $\mathcal{T}_{\phi,t}^{pruned}$.
@@ IV-B The $\log$-transform of constraints
The $\log$-transform is motivated by the desire to linearize the constraints related to the robustness computation in weight synthesis problems. Prior work [24, 22] demonstrated that constraints of the form $\rho(\sigma,\phi)>0$ can be encoded as a series of mixed-integer linear constraints. In the case of the weighted robustness, however, the temporal and logical operators multiply their operands’ robustness values by the associated weights (see Eq. (1)). Encoding these operations directly, following [12, 22], introduces multi-linearity whenever weights are treated as decision variables.
To address this, we apply a logarithmic transformation to both sides of mixed multi-linear integer constraints and convert the products into sums, thus linearizing them with respect to the weights. While effective in linearizing the product terms, the logarithm is only defined over the positive domain. In its most naive form, this requires both sides of every constraint to be positive. In other words, this transformation is valid only for signals that satisfy all predicates at all time instances that are used in the robustness computation (therefore, in constraints), which is not realistic. Before introducing our method for extending the transform to arbitrary signals (independent of satisfaction or violation), we first define the $\log$-transform of the robustness functions in Eq. (1) under the assumption that all predicate values are strictly positive. The $\log$-transform of the robustness of signal $\sigma$ at time $t$ is defined recursively as:
@@
We can still convert the right-hand sides of Eqs. (4) and (5) into a series of mixed-integer linear constraints using [22]. Note that we do not define the $\log$-transform for the robustness computation of negation. The restriction arises due to the domain in which the logarithm is defined: for a formula $\phi$, if $\log(r(\sigma,\phi))$ is defined, then $\log(r(\sigma,\lnot\phi))$ is not, and vice versa. In consequence, negation is not allowed in the formulas, which restricts our attention to formulas in PNF only. Negation in front of predicates is not an issue since it can be carried inside predicates by reverting the inequality.
When applying the $\log$-transform to constraints in Problem 1, $\log(h_{\mu}(\sigma(t))$ are computed apriori. The decision variables are $\log(w_{i})$. We perform a change of variables $v_{i}=\log(w_{i})$ during optimization, leaving the constraints $\log$-free. We can recover the original weights via $w_{i}=\exp(v_{i})$ afterward. Similarly, the robustness of the weighted formula can be recovered as $r(\sigma,\phi,t)=\exp(\log r(\sigma,\phi_{v},t))$ directly from the transformed constraints, without recomputing it from scratch. Next, we show that applying $\log$-transformation preserves the maximizer of Problem 1.
Theorem 2.
@@
Proof.
The logarithm is a strictly increasing function, and therefore preserves the ordering of all $\min$ and $\max$ operations in the formula. That is, for any set of $a_{1},\ldots a_{n}\in\mathbb{R}_{>0}$, we have $\log(\min(a_{1},\ldots,a_{n}))=\min(\log(a_{1}),\ldots,\log(a_{n}))$, and similarly for $\max$. Hence, while the numerical values of each operation in the Robustness Computation Tree change, the optimizers remain unchanged. ∎
@@
For signals whose robustness computation contains only negative robustness values, e.g., signals which violate all predicates of a formula at all time instances needed for the robustness computation, we cannot use $\log$-transform directly. However, because all these values lie in the negative quadrant, we can still compare them using the logarithm of their absolute values. Therefore, we can separate the sign from the magnitude, take the logarithm of the absolute value and multiply it with negative before performing the min/max operations, obtaining equations such as $\min_{i\in[1,2]}(-(\log(w_{i})+\log(|r(\sigma,\phi_{i},t)|))$. Equivalently, this can be handled by reversing the operator: for example, taking $\max$ instead of $\min$ when the values are negative, and removing the negative sign.
For signals whose robustness computation contains a mix of positive and negative values, we leverage the structural pruning algorithm to unify the signs in the computation. Applying structural pruning to a signal’s robustness computation allows us to remove any subtree that results in a robustness value whose sign is different than the sign of the robustness value at the root node. Pruning ensures that all remaining constraints have robustness values with the same sign, allowing the $\log$-transform to be applied safely to either all-positive robustness values or all-negative robustness values. Similar to the only-positive case, we show that applying these two procedures sequentially in a learning from human feedback problem does not change the maximizer of the problem.
Theorem 3.
@@ V Experiments
To evaluate the performance of the proposed method, we focus on two experiments: safe preference learning for robot navigation tasks, and interpretable learning-to-rank using real-life Formula 1 data.
@@ V-A Safe Preference Learning for Robot Navigation
where the state $q[t]=\begin{bmatrix}x&v_{x}&y&v_{y}\end{bmatrix}^{T}$ represents robot’s current position and velocity. We begin by randomly sampling weight valuations ten times to generate a trajectory for each of the WSTL formulas using PyTelo [25], and form five trajectory pairs. From these, we construct three preference datasets: PD1, the original set of preferences; PD2, obtained by flipping the preference of a single pair in PD1; PD3, obtained by inverting all preferences in PD1. For each preference dataset, we solve the MILP problem to learn the weight valuations that best represent the given preferences, yielding formulas $\phi_{w}^{\text{PD}i}$. Finally, we synthesize trajectories that will maximize the weighted robustness of the corresponding $\phi_{w}^{\text{PD}i}$. Figure 2 shows the synthesized trajectories. The results demonstrate that our method is sensitive to even small preference changes and reflects this diversity by producing distinct task executions.
@@ V-B Learning-to-rank
For the learning-to-rank example, we consider a real-world motorsport setting: Formula 1. Formula 1 is a wheel-to-wheel race where the goal is to finish in the highest position after a fixed number of laps. Drivers are ranked based on their finishing times, and each team, consisting of two drivers racing, heavily relies on on-track data to determine their race strategies, such as pit stops, tire compound selection, overtake plans, etc. Our goal is to learn a WSTL formula that captures the important aspects of racing on a specified track using previous years’ race data and ranking results. The learned WSTL formula represents an ideal race performance.
@@
A good race performance can be characterized by: starting from a competitive grid position, finishing the race, recording consistently fast lap times, outperforming (and overtaking) the car ahead, executing as few as necessary efficient pit stops, and avoiding on-track issues that affect the lap time, such as collisions and penalties [26]. To encode these principles in STL formalism, we consider several per-lap signals obtained via FastF1, a Python library for analyzing Formula 1 data [27]. Let $N$ be the number of laps run in a race. The percentage of laps completed by a car is $r_{lap}\in[0,1]^{N}$ . If a car retires early or is lapped, $r_{lap}<1$. The time-in-pit $t_{pit}\in\mathbb{R}^{N}$ records the duration of a pit stop on each lap, and is computed from the car’s entry and exit times in the pit lane (zero if no pit stop occurs). The fuel-and-pit-corrected lap times $\bar{t}_{lap}\in\mathbb{R}^{N}$ adjust the raw lap times for both pit-stop duration and the reduction in car mass due to fuel consumption, and computed for each lap $k\in\{1,\ldots,N\}$ as $\bar{t}_{lap}(k)=t_{raw}(k)-(m_{full}-c_{kg/lap}k)s_{s/kg}-t_{pit}(k)$ where $m_{full}=100$ represents the mass of the fully fueled car, $c_{kg/lap}$ represents the fuel use per each lap, and $s_{s/kg}=0.033$ represents mass sensitivity of lap time [28]. The position $p$ records the current race ranking of the car. The delta-to-lead $t_{\Delta lead}$ records the lap time differences between the ego and the car in front. Finally, the safety car presence $b_{SC}\in\{-1,1\}^{N}$ is set to $1$ during the Safety Car and Yellow Flag periods (i.e., when cars reduce speed) and 1-1 otherwise.
@@
Using these signals, we define the formula characterizing the performance as $\phi=\phi_{complete}\land\phi_{grid}\land\phi_{\Delta lead}\land\phi_{lap}\land\phi_{pit}$, where $\phi_{complete}=\Diamond(r_{lap}\geq 0)$ represents the effect of finishing the race, $\phi_{grid}=\square_{[0,1]}(p\leq 20)$ represents the effect of the starting position, $\phi_{\Delta lead}=\square(t_{\Delta lead}\leq 20s)$ represents the effect of competitiveness, $\phi_{lap}=\square(\lnot b_{SC}\implies\bar{t}_{lap}\leq 150s)$ represents the effect of the lap times excluding the reduce-speed periods, and $\phi_{pit}=\Diamond(t_{pit}>0\implies t_{pit}\leq 30s)$ represents the effect of having efficient pit stops.
@@
Figure 3: The evolution of final standing predictions and the prediction accuracy over the laps at the 2025 Monza Grand Prix. The right block shows the correct standing when DNFs are included and excluded. Each team is represented with a separate color, and drivers from the same team are represented with solid and dashed lines. For readability purposes, driver abbreviations are kept instead of enumerations.
@@
$\text{acc}((\sigma_{1},\ldots,\sigma_{P}))=\frac{\text{number of discordant pairs}}{{P\choose 2}}$.
@@
Experiment details: To ensure consistent racing conditions over the years, such as weather and safety car appearances, we focus on the Monza Grand Prix. Since there are no major regulation changes in recent years, we use data from 2021, 2022, and 2024 for training, and the 2025 race for testing. The 2023 Monza Grand Prix is excluded due to a false start that shortened the race to $51$ laps instead of the usual $53$. For both training and test data, we scale all signals at each year to $[0,1]$ by min-max normalization, to avoid biases due to the order-of-magnitude difference across channels. Because real races involve unpredictable events such as collisions and technical failures, we perform two sets of experiments: one excluding Did Not Finish (DNF) and Did Not Start (DNS) cars and one including them.
@@
For both cases, we warm start the MILP solver by selecting the best-accuracy solution among $S=10000$ uniform random weight samples (cf. Random Sampling (RS) method of [10]). We then solve the MILP problem with a four-hour time limit, and report the best solution found. As a baseline comparison, we repeat the RS method $1000$ times with $S_{RS}=1000$ weight samples, and report the mean and standard deviations of accuracies. Table I shows the results on training and test sets. It is important to note that even though a global optimal is not guaranteed when the solver is time-limited, the MILP procedure improves over mean RS performance, which shows almost $10\%$ variability on the test accuracy. Also, when compared to its warm-started accuracy, our method raises the accuracy by up to $5\%$. Moreover, the learned weights generalize to future seasons with different cars, different teams, and different drivers. This highlights the fact that, although driver skills and car performances are at the core of the race success, the proposed method can capture driver-and car-agnostic performance points.

### v2

@@ III Problem Statement
Learning from human feedback in autonomous systems encompasses problems such as learning from preferences, demonstrations, or rankings. Classical approaches include inverse reinforcement learning and behavioral cloning, which directly infer policies or reward structures from data [21, 22]. Both methods typically require a reinforcement learning-style framework. In contrast, we propose to learn a utility function that captures the preferences independent of a specific agent or environment. Ideally, this utility function serves as a unifying representation of human intent, and whenever trajectories consistent with these preferences are required, they can be synthesized via downstream trajectory generation methods. To restrict the search space of possible utility functions, we focus on a structured family of functions defined by the safety requirements or task specifications. Particularly, we leverage the expressive power of WSTL to describe specifications and the weighted robustness. By parameterizing weights in the quantitative semantics, we enable different weight valuations to encode potential instantiations of the utility function. The learned WSTL formula can be seamlessly integrated to control synthesis frameworks such as model predictive control (MPC), to generate safe and personalized trajectories. Moreover, as weight valuations represent the importance of subformulas or time instances, the result of the learning framework is interpretable, in contrast to using neural network-like structures for learning.
In other words, solving learning-from-human-feedback problems via WSTL requires a task description written in STL and a user dataset, and searches for weighted quantitative semantics that maximize the problem’s objective. Formally, we define the problem as follows:
Problem 1 (Learning from Human Feedback (LHF))
@@
$\begin{array}[]{ccc}w^{*}\in&\max\limits_{w\in\mathbb{R}_{>0}^{n}}&f(\phi_{w},\mathcal{D})\\ &\text{ s.t. }&\text{dataset constraints}\end{array}.$ (3)
The form of the objective function and dataset constraints depends on the learning setup and the available data. For safe preference learning, the user dataset is given as $\mathcal{PR}=\{(\sigma_{i_{1}},\sigma_{i_{2}},l_{i})\}_{i=1}^{P}$, where $\sigma_{i_{1}}$ and $\sigma_{i_{2}}$ are compared and $l_{i}\in\{1,-1\}$ denotes its label: $l_{i}=1$, when $\sigma_{i_{1}}$ is preferred over $\sigma_{i_{2}}$ and $l_{i}=-1$ otherwise. The goal is to maximize the number of correctly ordered pairs, leading to the objective function of the form $f(\phi_{w},\mathcal{PR})=\sum_{i=1}^{P}\mathbb{1}_{\big(l_{i}\Delta_{i}(w)>0\big)}$, and constraints $\Delta_{i}(w)=r(\sigma_{i_{1}},\phi_{w})-r(\sigma_{i_{2}},\phi_{w})$.
For learning to rank, the dataset is an ordered set $\mathcal{R}=\{\sigma_{1}\succeq\sigma_{2}\succeq\ldots\succeq\sigma_{P}\}$. This can be reformulated as a pairwise preference dataset $\mathcal{PR}=\{(\sigma_{i},\sigma_{j},1)\}_{i=1,j=i+1}^{P-1,P}$, reducing the problem to safe preference learning with the same objective function and $\binom{P}{2}$ constraints. Finally, in learning from demonstrations, we are given a set of “desired” trajectories $\mathcal{D}=\{\sigma_{i}\}_{i=1}^{D}$. In this setup, we can use $f(\phi_{w},\mathcal{D})=\sum_{i=1}^{D}r(\sigma_{i},\phi_{w})$ as the objective value, which aims to assign high weighted robustness values to desired trajectories. When dataset constraints are chosen as above, we can restrict the search domain for weights from strictly positive reals to a bounded region $a_{1}<w_{i}\leq a_{2},\;i\in[1,n]$, where $0\leq a_{1}<a_{2}$. The proof of feasibility equivalence can be found in Th. 2 [10]. In all cases, the cost function may include regularization terms in the form of $f(\phi_{w},\mathcal{PR})+\lambda|w|_{p}$, where $|w|_{p}$ denotes the $p$-norm of the vector of all weights, with a slight abuse of notation.
The computational complexity of WSTL synthesis problems grows with the complexity of the formula. In general, computing the weighted robustness for a signal involves evaluating predicate values at multiple time instances, multiplying them with corresponding weight valuations, and applying nested min/max operations, following the WSTL semantics. Since weights appear multiplicatively, optimizing over them results in a mixed integer program (MIP) with multi-linear constraints, where integers appear as in [14]. As such, weight synthesis problems are solved with heuristic approaches such as gradient descent [23] or sampling [10] in the literature.
@@ IV Reduction to MILP
In this paper, we propose two procedures to reduce the problem size and linearize the constraints, making the problem easier to solve. The first procedure, called structural pruning, systematically examines how values in the Robustness Computation Tree are propagated to the root node to determine the robustness value of a signal with respect to a formula. Based on this observation, we can prune the branches that cannot affect the robustness value regardless of the weight valuations assigned to them. The second procedure, called the $\log$-transform, leverages the product identity of the logarithm function to convert constraints in the multiplication form into constraints in the summation form. Since the logarithm is only defined for positive values, a naive approach requires that all subformulas (subtrees in RCT) must have positive robustness for the transformation to be valid, which is a restrictive requirement in practice. To address this, we combine the $\log$-transform with structural pruning, and ensure that only contributing parts with the same robustness value sign of the formula are considered in the robustness computation, therefore making all variables of the constraints that encode this computation have the same sign. Together, these two procedures allow the problem to be reformulated as an MILP with a potentially reduced number of constraints.
Before describing the procedures in detail, we note that all formulas are assumed to be in positive normal form (PNF), that is, negations appear only in front of predicates. While this assumption is not restrictive, as negations can be pushed to the predicate level, we discuss the rationale later.
@@ IV-A Structural Pruning
The key insight of structural pruning is that if the overall robustness of a signal is positive (indicating satisfaction), then any subformula with negative or zero robustness values must be “absorbed” by one of its parent operations in the Robustness Computation Tree. Consequently, when fitting weights to this tree, we know that weights connecting nodes with negative or zero robustness value to their parents cannot change the robustness. This allows us to systematically prune parts of the tree whose values are not actively contributing to robustness.
@@
Function Prune($\sigma,\phi,t,s_{\phi}$):
 if Root node of $\mathcal{T}_{\phi,t}$ is an Operator then
    for $(\phi_{i},t^{\prime})\in$ Children of the Root do
       if $\mathrm{sign}(\rho(\sigma,\phi_{i},t^{\prime}))=s_{\phi}$ then
          $\mathcal{T}_{\phi_{i},t}\leftarrow$ Prune($\sigma,\phi_{i},t^{\prime},s_{\phi_{i}}$)
       else
          $\mathcal{T}_{\phi,t}\leftarrow\mathcal{T}_{\phi,t}\setminus\mathcal{T}_{\phi_{i},t^{\prime}}$
 else
    return $\mathcal{T}_{\phi,t}$
-0.2em
@@
Example 1 (continued)
Consider $\phi$, and $\sigma=[5,6,7,-1,4,2]$. Figure 1(c) shows the pruned RCT, $\mathcal{T}_{\phi,t}^{pruned}$.
@@ IV-B The $\log$-transform of constraints
The $\log$-transform is motivated by the desire to linearize the constraints related to the robustness computation in weight synthesis problems. Prior work [24, 14] demonstrates that constraints of the form $\rho(\sigma,\phi)>0$ can be encoded as a series of mixed-integer linear constraints. In the case of the weighted robustness, however, the temporal and logical operators multiply their operands’ robustness values by the associated weights (see Eq. (II-B)). Encoding these operations directly, following [12, 14], introduces multi-linearity whenever weights are treated as decision variables.
To address this, we apply a logarithmic transformation to both sides of mixed multi-linear integer constraints and convert the products into sums, thus linearizing them with respect to the weights. While effective in linearizing the product terms, the logarithm is only defined over the positive domain. In its most naive form, this requires both sides of every constraint to be positive. In other words, this transformation is valid only for signals that satisfy all predicates at all time instances that are used in the robustness computation (therefore, in constraints), which is not realistic. Before introducing our method for extending the transform to arbitrary signals (independent of satisfaction or violation), we first define the $\log$-transform of the robustness functions in Eq. (II-B) under the assumption that all predicate values are strictly positive. The $\log$-transform of the robustness of signal $\sigma$ at time $t$ is defined recursively as:
@@
We can still convert the right-hand sides of Eqs. (4) and (5) into a series of mixed-integer linear constraints using [14]. Note that we do not define the $\log$-transform for the robustness computation of negation. The restriction arises due to the domain in which the logarithm is defined: for a formula $\phi$, if $\log(r(\sigma,\phi))$ is defined, then $\log(r(\sigma,\lnot\phi))$ is not, and vice versa. In consequence, negation is not allowed in the formulas, which restricts our attention to formulas in PNF only. Negation in front of predicates is not an issue since it can be carried inside predicates by reversing the inequality.
When applying the $\log$-transform to constraints in Problem 1, $\log(h_{\mu}(\sigma(t)))$ are computed a priori. The decision variables are $\log(w_{i})$. We perform a change of variables $v_{i}=\log(w_{i})$ during optimization, leaving the constraints $\log$-free. We can recover the original weights via $w_{i}=\exp(v_{i})$ afterward. Similarly, the robustness of the weighted formula can be recovered as $r(\sigma,\phi,t)=\exp(\log r(\sigma,\phi_{v},t))$ directly from the transformed constraints, without recomputing it from the original constraints. Next, we show that applying $\log$-transformation preserves the maximizer of Problem 1.
Theorem 2
@@
Proof:
The logarithm is a strictly increasing function, and therefore preserves the ordering of all $\min$ and $\max$ operations in the formula. That is, for any set of $a_{1},\ldots,a_{n}\in\mathbb{R}_{>0}$, we have $\log(\min(a_{1},\ldots,a_{n}))=\min(\log(a_{1}),\ldots,\log(a_{n}))$, and similarly for $\max$. Hence, while the numerical values of each operation in the Robustness Computation Tree change, the optimizers remain unchanged. ∎
@@
For signals whose robustness computation contains only negative robustness values, e.g., signals which violate all predicates of a formula at all time instances needed for the robustness computation, we cannot use $\log$-transform directly. However, because all these values lie in the negative quadrant, we can still compare them using the logarithm of their absolute values. Therefore, we can separate the sign from the magnitude, take the logarithm of the absolute value and multiply it with negative before performing the min/max operations, obtaining equations such as $\min_{i\in\{1,2\}}(-(\log(w_{i})+\log(|r(\sigma,\phi_{i},t)|))$. Equivalently, this can be handled by reversing the operator: for example, taking $\max$ instead of $\min$ when the values are negative, and removing the negative sign.
For signals whose robustness computation contains a mix of positive and negative values, we leverage the structural pruning algorithm to unify the signs in the computation. Applying structural pruning to a signal’s robustness computation allows us to remove any subtree that results in a robustness value whose sign is different than the sign of the robustness value at the root node. Therefore, pruning is a required step for the $\log$-transform to be applicable, as it ensures that all remaining constraints have robustness values with the same sign. As in the all-positive case, we show that applying these two procedures sequentially in a learning-from-human-feedback problem does not change the maximizer of the problem.
Theorem 3
@@ V Experiments
To evaluate the performance of the proposed method, we focus on two experiments: safe preference learning for robot navigation tasks, and interpretable learning to rank using real-life Formula 1 data. The implementation details are available at https://github.com/ruyakrgl/human2wstl.git.
@@ V-A Safe Preference Learning for Robot Navigation
where the state $q[t]=\begin{bmatrix}x&v_{x}&y&v_{y}\end{bmatrix}^{T}$ represents robot’s current position and velocity. We begin by randomly sampling weight valuations ten times to generate a trajectory for each of the WSTL formulas using PyTelo [25], and form five trajectory pairs. From these, we construct three preference datasets: PD1, the original set of preferences; PD2, obtained by flipping the preference of a single pair in PD1; PD3, obtained by inverting all preferences in PD1. For each preference dataset, we solve the MILP problem to learn the weight valuations that best represent the given preferences, yielding formulas $\phi_{w}^{\text{PD}i}$. Finally, we synthesize trajectories that will maximize the weighted robustness of the corresponding $\phi_{w}^{\text{PD}i}$. Figure 2 shows the synthesized trajectories. The results demonstrate that our method is responsive to even small preference changes and captures this diversity by producing distinct task executions.
@@ V-B Learning to Rank
For the learning-to-rank example, we consider a real-world motorsport setting: Formula 1. Formula 1 is a wheel-to-wheel racing championship where the goal is to finish in the highest position after a fixed number of laps. Drivers are ranked based on their finishing times, and each team, with two drivers per race, heavily relies on on-track data to determine their race strategies, such as pit stops, tire compound selection, overtake plans, etc. Our goal is to learn a WSTL formula that captures the important aspects of racing on a specified track using previous years’ race data and ranking results. The learned WSTL formula represents an ideal race performance. Note that, although this method is introduced as safe learning, this example shows that the approach can be extended to any kind of task execution guarantee written in STL.
@@
A good race performance can be characterized by: starting from a competitive grid position, finishing the race, recording consistently fast lap times, outperforming (and overtaking) the car ahead, executing as few efficient pit stops as necessary, and avoiding on-track issues that affect the lap time, such as collisions and penalties [26]. To encode these principles in STL formalism, we consider several per-lap signals obtained via FastF1, a Python library for analyzing Formula 1 data [27]. Let $N$ be the number of laps run in a race. The percentage of laps completed by a car is $r_{lap}\in[0,1]^{N}$. If a car retires early or is lapped, $r_{lap}(k)<1$, for all laps $k\in\{1,\ldots,N\}$. The time-in-pit $t_{pit}\in\mathbb{R}^{N}$ records the duration of a pit stop on each lap, and is computed from the car’s entry and exit times in the pit lane (zero if no pit stop occurs). The fuel-and-pit-corrected lap times $\bar{t}_{lap}\in\mathbb{R}^{N}$ adjust the raw lap times for both pit-stop duration and the reduction in car mass due to fuel consumption, and is computed for each lap $k\in\{1,\ldots,N\}$ as $\bar{t}_{lap}(k)=t_{raw}(k)-(m_{full}-c_{kg/lap}k)s_{s/kg}-t_{pit}(k)$ where $m_{full}=100$ represents the mass of the fully fueled car, $c_{kg/lap}$ represents the fuel use per each lap, and $s_{s/kg}=0.033$ represents mass sensitivity of lap time [28]. The position $p$ records the current race ranking of the car. The delta-to-lead $t_{\Delta lead}$ records the lap time differences between the ego and the car in front. Finally, the safety car presence $b_{SC}\in\{-1,1\}^{N}$ is set to $1$ during the Safety Car and Yellow Flag periods (i.e., when cars reduce speed) and 1-1 otherwise.
@@
Using these signals, we define the formula characterizing the performance as $\phi=\phi_{progress}\land\phi_{grid}\land\phi_{\Delta lead}\land\phi_{lap}\land\phi_{pit}$, where $\phi_{progress}=\Diamond(r_{lap}\geq 0)$ represents the effect of the progress, $\phi_{grid}=\square_{[0,1]}(p\leq 20)$ represents the effect of the starting position, $\phi_{\Delta lead}=\square(\lnot b_{SC}\implies t_{\Delta lead}\leq 30s)$ represents the effect of competitiveness, $\phi_{lap}=\square(\lnot b_{SC}\implies\bar{t}_{lap}\leq 150s)$ represents the effect of the lap times excluding the reduce-speed periods, and $\phi_{pit}=\square(t_{pit}>0\implies t_{pit}\leq 30s)$ represents the effect of having efficient pit stops. To compare and analyze the contribution of the formula, we also use $\tilde{\phi}=\phi_{progress}\land\phi_{grid}\land\phi_{\Delta lead}\land\phi_{pit}$. Note that $\tilde{\phi}$ is the same except $\phi_{lap}$ is not in the specification.
@@
Figure 3: The evolution of the final standing predictions at the 2025 Monza Grand Prix when (a) DNF and DNS are excluded, and (b) DNF and DNS are included. For each subfigure, the plot on the left shows the predictions after each lap, and the block on the right shows the true final standings. Each team is represented with a separate color, and drivers from the same team are represented with solid and dashed lines. For readability, driver abbreviations are used instead of enumerations.
@@
$\text{acc}((\sigma_{1},\ldots,\sigma_{P}))=1-\frac{\text{number of discordant pairs}}{{\binom{P}{2}}}$.
@@
Experiment details: To ensure consistent racing conditions over the years, such as weather and safety car appearances, we focus on the Monza Grand Prix. Since there are no major regulation changes in recent years, we use data from 2021, 2022, and 2024 for training, and the 2025 race for testing. The 2023 Monza Grand Prix is excluded due to a false start that shortened the race to $51$ laps instead of the usual $53$. For both training and test data, we scale all signals and thresholds at each year to $[0,1]$ by min-max normalization, to avoid biases due to the order-of-magnitude difference across channels. Because real races involve unpredictable events such as collisions and technical failures, we perform two sets of experiments: one excluding Did Not Finish (DNF) and Did Not Start (DNS) cars and one including them.
@@
For both cases, we warm-start the MILP solver by selecting the best-accuracy solution among $S=10000$ uniform random weight samples (cf. Random Sampling (RS) method of [10]). We set the cost function as defined in Sec. III with the regularization term $\lambda=0.2$. We then solve the MILP problem with a four-hour time limit, using Gurobi on a computer with 64 AMD Ryzen Threadripper PRO 5975WX 32-Cores and report the best solution found. As a baseline comparison, we repeat the RS method $1000$ times with $S_{RS}=10000$ weight samples, and report the mean and standard deviations of accuracies. RS is shown to be comparable to Gradient-Descent methods in terms of accuracy, and more efficient in terms of computation time [10]. We do not compare with the naive version (without pruning and log-transform) as MINLPs are highly costly, and the proposed method outperforms the MINLP problem by construction, i.e., a simpler and smaller problem. Table I shows the results on training and test sets for both DNF options and both formulas. It is important to note that even though a global optimal is not guaranteed when the solver is time-limited, the MILP procedure improves over mean RS performance on the training set and shows promising performance on the test set. Also, when compared to its warm-started accuracy, our method raises the accuracy by up to $7\%$. Moreover, when $\phi$ is used, the learned weights generalize to future seasons with different cars, different teams, and different drivers. This highlights the fact that, although driver skills and car performances are at the core of the race success, the proposed method can capture driver- and car-agnostic performance points. We do not observe the same performance with $\tilde{\phi}$, specifically, when DNF/DNS are included, the formula overfits to previous years. Options to fix are increasing the regularization constant or choosing a more representative formula.